OpenStudy (anonymous):

evaluate lim x->0 (1/x^3) multiplied by antiderivative of ln(1+t^2)from the bounds of x to 0

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0}\frac{ 1 }{ x^3 }\int\limits_{0}^{x} \ln(1+t^2)dt\]

OpenStudy (anonymous):

u multiply in the 1/x^3 to make it more simple

OpenStudy (anonymous):

then i don't know why  is in indeterminate form im so tired